OpenStudy (anonymous):

How do I factor e^2x + 2e^x - 8 ?

OpenStudy (anonymous):

Find the real solutions of e2x + 2ex 8 = 0

OpenStudy (anonymous):

same way you'd factor a polynomial: all techniques apply since e^x * e^x = e^2x so its: (e^x + 4)(e^x - 2)

OpenStudy (anonymous):

@lukepuller \(e^{2x}=(e^x)^2\) so let \(u=e^x\) and we get:$$e^{2x}+2e^x-8=u^2+2u-8$$can you factor \(u^2+2u-8\)? once you do, plug back in \(u=e^x\)

OpenStudy (anonymous):

thanks!